Suppose​ 36% of​ movie-goers say that horror is their favorite movie genre. If a random

sample of 120​ movie-goers is​ taken, what are the mean and standard deviation of the

number of​ movie-goers that prefer​ horror?
Answer #1

Solution :

Given that,

p = 36%=0.36

q = 1 - p =1-0.36=0.64

n = 120

Using binomial distribution,

mean = = n * p = 120*0.36=43.2

standard deviation   = n * p * q = 120*0.36*0.64=5.2581